OpenStudy (anonymous):

Solve and graph the inequality. 4/5x + 5 < -3

OpenStudy (anonymous):

x=-10

OpenStudy (anonymous):

X<-10

OpenStudy (anonymous):

\[\frac{ 4 }{ 5x+5 }<-3\] or \[\frac{ 5x+5 }{ 4 }>-3\] multiply by 4 5x+5>-12 5x>-12-5 5x>-17 \[x>\frac{ -17 }{ 5 }\] x>-3.4 now you can make the graph.